OpenStudy (anonymous):

Rationalize the denominator. The answer should be expressed in simplified form.

OpenStudy (anonymous):

\[\frac{ \sqrt{x}+3 }{ \sqrt{x}-3 }\]

OpenStudy (anonymous):

you need to use the inverse function to simplified the expression

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

\[\times \sqrt{x}+3 \] top and bottom

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

\[\frac{ x +9}{ x-3 }\]?

OpenStudy (shiraz14):

Multiply by (√x+3)/(√x+3).

OpenStudy (anonymous):

So it is... x-9/x-3?

OpenStudy (anonymous):

Or x-9/x-9?

OpenStudy (shiraz14):

No, it would be (x+6√x+9)/(x-9).

OpenStudy (anonymous):

Oh ok, Is that is or there's more?

Miracrown (miracrown):

we want to get rid of the radicals from the denominator it makes sense with the play on words, if our denominator has a radical in it then its crazy and not very rational. it radical! so we want to rationalize it and that means getting rid of the radical

OpenStudy (shiraz14):

Well it can be further simplified into the following: 1 + 6(√x-3)/(x-9)

Miracrown (miracrown):

@cookiibabii93 I'm trying to help you understand exactly what you need to do so in future you can apply that. :)

OpenStudy (anonymous):

Oh ok, thanks @Miracrown & @shiraz14

OpenStudy (shiraz14):

@Miracrown is correct.

Miracrown (miracrown):

No worries. :-]